# Point2Primitive: CAD Reconstruction from Point Cloud by Direct Primitive Prediction

### v2

@@ Abstract
Recovering CAD models from point clouds, especially the sketch-extrusion process, can be seen as the process of rebuilding the topology and extrusion primitives. Previous methods utilize implicit fields for sketch representation, leading to shape reconstruction of curved edges. In this paper, we proposed a CAD reconstruction network that produces editable CAD models from input point clouds (Point2Primitive) by directly predicting every element of the extrusion primitives. Point2Primitive can directly detect and predict sketch curves (type and parameter) from point clouds based on an improved transformer. The sketch curve parameters are formulated as position queries and optimized in an autoregressive way, leading to high parameter accuracy. The topology is rebuilt by extrusion segmentation, and each extrusion parameter (sketch and extrusion operation) is recovered by combining the predicted curves and the computed extrusion operation. Extensive experiments demonstrate that our method is superior in primitive prediction accuracy and CAD reconstruction. The reconstructed shapes are of high geometrical fidelity.
Figure 1: CAD reconstruction pipeline including (a) The Point2Primitive network first segmented the input point cloud into clusters belonging to the same extrusion and corresponding barrel points, and the curves of each extrusion primitive are directly learned and predicted. CAD reconstruction can be achieved by combining the reconstructed extrusion sketch and extrusion operations computed from the barrel points, which is easy to edit in downstream tasks. (b) The illustrations of the sketch-extrude process, extrusion segmentation, and extrusion primitives. The topology describes how extrusion primitives are interrelated.
@@ 1 Introduction
CAD reconstruction from point clouds is a longstanding objective [10]. In CAD, reverse engineering encompasses techniques that extract CAD models from existing products by transforming unstructured data (e.g. point clouds, meshes, or printed schematics) into structured representations of 2D or 3D geometries. The sketch-and-extrude process is one of the most fundamental CAD modeling commands, offering a highly intuitive and versatile way to generate a broad array of shapes, as demonstrated in [36], and investigating the reverse of this process is valuable to the community.
However, the unordered and unstructured nature of point clouds poses significant challenges for high-level manipulation and efficient editing of their underlying geometries [27]. Some methods try to solve this problem and reconstruct the shapes through optimization methods [19, 1, 30, 15, 18, 26]. The recent availability of significant boundary representation (B-Rep) datasets [36, 9, 31] has introduced data-driven frameworks to reverse engineering in the CAD world. Numerous neural networks have been developed to segment point clouds into clusters to represent distinct surface regions [13, 32, 11, 21]. These segmented points are then fitted to surface primitives (either parametric or freeform), allowing the topological assembly of these primitives to reconstruct the CAD models. In contrast, we focus on recovering the CAD models that describe the sketch-and-extrude modeling process. To this end, we employ a point segmentation network to recover topological structures (extrusion primitives) by segmenting the point clouds into clusters associated with individual extrusion primitives. By predicting the curves and extrusion operation of each extrusion primitive, the CAD models can be reconstructed.
Specifically, the sketch serves as the foundation of parametric CAD modeling, with the quality of the final CAD model depending critically on the precision of its sketches [5]. In the parametric CAD workflow, close curves constitute the sketch, and 3D models are then constructed by extruding the sketches and grouping these extrusions into structured assemblies. Recently, implicit representations of 3D shapes [35, 23], have inspired new approaches that employ SDFs to represent sketches, enabling sketch and CAD reconstruction through SDF encodings [34, 29, 14]. However, decoding SDF embeddings into sketches typically results in CAD models with unintended curved edges and the reconstructed sketch is not easy to be directly edited. Inspired by large language models (LLMs) [4], recent works [37, 39, 22] propose to generate CAD modeling sequences using language-based models. However, our findings indicate that LLMs struggle to implicitly learn hierarchical representations of topology (e.g. CSG structures) [41] using only reconstruction loss, such as chamfer distance and primitive loss, which can result in limited generalization. SkexGen [38] strives to address this challenge by introducing separate sequence encoders for modeling topology and geometry independently to facilitate related design generation. In response, we propose representing sketches through their fundamental curves and explicitly recovering their extrusion primitives [41] by decomposing input points into distinct extrusion sets.
This paper presents a novel approach for reconstructing sketch-and-extrude CAD models using an end-to-end point-to-primitive (Point2Primitive) network. Specifically, our method first segments the input point cloud into distinct extrusion primitives and corresponding barrel/base points. Each extrusion sketch is then predicted by directly predicting the type and parameter of its curves through an improved transformer network. By combining the predicted curves and the extrusion operation calculated from the barrel points, the CAD models can be rebuilt. Using curves as sketch representation, our approach reconstructs fundamental shapes in sketches (such as lines, arcs, and circles), resulting in accurate, geometry-preserving, and easily editable 3D CAD models. Our primary contributions are summarized as follows:
  • We propose a CAD reconstruction method based on the Point2Primitive network, which leverages an enhanced transformer to predict sketch and PointNet++ for extrusion segmentation. By representing the sketch through its fundamental curves rather than using SDF, our approach achieves higher geometric fidelity.
  • We introduce an improved transformer decoder along with a sketch hierarchy representation. The curve parameters are encoded as position embeddings to guide the attention weight in locating the target. The parameters are directly predicted and optimized in an auto-regressive way, leading to high accuracy.
  • We present a complete data generation pipeline that can produce extrusion segmentation labels, barrel/base segmentation labels, and sketch primitives from the CAD modeling sequence saved in JSON files, which can be used for future research.
  • We conduct a comprehensive experiment with several CAD reconstruction results to analyze the proposed method thoroughly. Also, we reproduced the data generation methods used by other methods for comparison. An ablation study is conducted to fully investigate the proposed method.
@@ 2.1 CAD Reconstruction via Primitive Detection
Several learning-based approaches have also been proposed to fit geometric primitives to point clouds. These methods first segment the input points into clusters corresponding to the same surface primitives and fit the clustered points to the parametric surface primitives [13, 32, 11, 17] or free surfaces [21]. Except for surface primitives, UCSG-Net[7] and CSG-stump [28] learn the CSG parse tree that interrelates the predicted geometry primitives to reconstruct shapes. The reconstructed CAD models are in B-rep format, which is not as easy to edit as the sketch-and-extrude modeling procedure. In this paper, we explore the inverse of the sketch-and-extrude process and strive to recover the CAD modeling procedure from the point cloud. Therefore, we segment the point clouds into clusters belonging to the same extrusion primitive and explicitly reconstruct every element of the extrusion primitives. Instead of fitting points to the parametric or free surface, we employ an improved transformer to learn and predict the curves directly.
@@ 2.2 Sketch-and-extrude CAD Reconstruction
Some new solutions use SDF to represent sketches following the recent implicit 3D shape representations [23, 35]. ExtrudeNet [27] and SECAD-Net [14] can learn implicit sketches and differentiable extrusions. Point2Cyl reconstructs 3D shapes through sketch regression supervised by the latent embeddings of its SDF. However, these methods utilize the implicit fields for sketch representation, leading to curved edges of the reconstructed shapes. Meanwhile, the reconstructed sketch is hard to be directly edited because it is represented by the SDF. Further transformation from SDF to sketch further brings in extra parameter errors. In this paper, we propose to represent a sketch using its curves and directly learn and predict the curve types and parameters from the point clouds in an object detection manner [33]. To this end, we propose the center-prior primitive definition (detailed in Section 3) and the improved transformer decoder (detailed in Section 4.4) to achieve precise and editable sketch reconstruction, which makes the final CAD reconstruction easy to be edited. Also, the most related work to us is the Point2Cyl, and we conduct an extra comparison as briefed in Section .
@@ 2.3 CAD Generation
Following Large Language Model (LLM) [4], [37] is the first to explore the CAD sequence generation of 3D shapes based on VAE [8]. [39] utilizes a vector quantized VAE to generate modeling sequences. [22] propose to generate CAD models step-by-step based on multimodal diffusion. It should be noted that the CSG structure [41] is broken when tokenizing the CAD modeling sequence. We find that the implicit representation learning of the topology is challenging for LLM if only supervised by geometry loss (such as Chamfer Distance) or command accuracy. SkexGen [38] propose to use two sequence encoders separately for topology and geometry learning. However, CAD generation methods leveraging VAE [8] structure and language model decoder still limit its performance on reconstruction tasks because of poor generalization. Also, language models struggle with numerical representation and mathematics [40], causing inaccuracies in command parameters and limited performance beyond the training dataset (detailed in Section 5.4). In this paper, we present a CAD reconstruction method based on the Point2Primitive network. We explicitly recover the topology through point segmentation and directly predict the sketch curves and extrusion operation. By explicit precise primitive parameter prediction, the CAD reconstructions are of high geometry fidelity and good generalization.
3 Preliminaries
In CAD terminology, the sketch consists of multiple closed curves $S_{i}$. There are three types of curves: line $\mathit{L}$, arc $\mathit{A}$, and circle $\mathit{R}$ in our sketch hierarchy. Therefore, the sketch $S_{i}=[c_{1}^{i},...,c_{N}^{i}]$, where $c_{j}^{i}\in(\mathit{L},\mathit{A},\mathit{R})$ for all $j\in[1,...,N]$, can be seen as a set of curves. The goal of predicting sketch is to train the parameters $\theta$ of the Point2Primitive network $f_{\theta}$ on the training pair $({p_{i}},\hat{S}_{i})$ to make a good approximator of the set prediction as follows:
$f_{\theta}\left(\{p_{k}\mid p_{k}\in P_{\mathbb{E}_{i}}\}\right)\approx\left\{c_{1}^{i},...,c_{j}^{i}\mid c_{j}^{i}\in(\mathit{L},\mathit{A},\mathit{C})\right\}$ (1)
where $P_{\mathbb{E}_{i}}\in R^{N_{E_{i}}\times{3}}$ is the points of the $i\mbox{-}th$ extrusion $\mathbb{E}_{i}$.
Curve Type Parameters
$L$ $(Line)$ $x_{m}$ $y_{m}$ $x_{1}$ $y_{1}$1 -1
$A$ $(Arc)$ $x_{m}$ $y_{m}$ $x_{1}$ $y_{1}$ $x_{2}$ $y_{2}$
$R$ $(Circle)$ $x_{c}$ $y_{c}$ $r$1 -1 -1
Table 1: Curve definition. We provide the types and vector representations of curves used in this work. Unused parameters in curve vector are padded with $-1$.
To work with the improved transformer decoder, we formulate the definition of curve parameter in a centra-prior way, as shown in Table 1. Each primitive parameter is primarily presented by its center.
Line is represented with its midpoint ($x_{m},y_{m}$) and another endpoint ($x_{1},y_{1}$). Arc is defined with its midpoint ($x_{m},y_{m}$), start points ($x_{1},y_{1}$) and endpoints ($x_{2},y_{2}$). The circle parameter consists of the center point ($x_{c},y_{c}$) and radius $r$. Each curve $C_{i}=(t_{i},\rho_{i})$ is defined by its primitive types $t_{i}\in{\mathbb{R}}$ and primitive parameters $\rho_{i}\in\mathbb{R}^{1\times{6}}$ with unused parameters are padded with $-1$. Therefore, the first two elements of the parameters $\rho_{i}$ can be encoded as position embeddings to guide the attention in the improved transformer decoder as briefed in Section 4.4
In the sketch-extrude process, a sketch $S_{i}$ is extruded by an operation $E_{i}$ to form an extrusion $\mathbb{E}_{i}=\left[S_{i},E_{i}\right]$, and multiple such extrusions collectively form a solid shape $\mathbb{S}=\{\mathbb{E}_{1},...,\mathbb{E}_{K}\}$. The shape $\mathbb{S}$, therefore, can be represented by the SDF defined by the collection of the extrusions. In our method, the $\{\mathbb{E}_{1},...,\mathbb{E}_{i}\}$ is reconstructed via point segmentation, while the sketch $S_{i}$ is predicted through an improved transformer network.
4 Point2Primtive
4.1 Overview
We propose Point2Primitive, a deep model for CAD reconstruction from point clouds. As illustrated in Figure 2, our method utilizes a point segmentation network to cluster points that belong to the same extrusion. Base/barrel points are also classified during extrusion segmentation. Then an improved transformer is introduced to predict the curve type and parameters of each extrusion. As defined in Equation 1, curve prediction can be seen as a set prediction problem. Therefore, we use the transformer network to solve the set-to-set problem and make some improvements to the transformer decoder, as shown in Figure 2. The transformer encoder is used to refine the barrel point features from the segmentation (Seg) head. The Seg head is detailed in Appendix 4.2, where the barrel point encoder is developed based on the PointCNN [16] with fixed grid points as representation points instead of random points.
@@ 4.2 CAD Dataset Generation
In this paper, we propose to generate the training dataset based on the Signed Distance Field (SDF). An SDF $SDF(p_{eval},b)$ is a continuous function that, for a given spatial eval point $p_{eval_{i}}$, outputs the eval point’s distance to the closest boundary defined by $b$, whose sign denotes whether the point is inside (negative) or outside (positive) of the watertight surface. The data generation pipeline can convert any CAD sequence saved in JSON files following the style of the Fusion360 Gallery dataset [36] to the training data pairs.
Given an extrusion point set $P_{\mathbb{E}_{i}}\in R^{N_{\mathbb{E}_{i}}\times{3}}$ and its sketch $S_{i}$. We first project the points $P_{\mathbb{E}_{i}}$ to the sketch plane as eval points according to the transform matrix $W_{i}$ defined by its extrusion axis $e_{i}$. Then, its SDF is calculated to the boundary defined by the sketch $S_{i}$. The batch SDF calculation will be detailed in the Appendix 5.1. The barrel label $b_{i}$ of each point can be defined as follows:
$b_{i}=\text{bool}(\text{SDF}(W_{i}\cdot{P_{\mathbb{E}_{i}}},S_{i})\leq\text{thresh})$ (2)
where $thresh$ controls the label precision and $\text{thresh}=0.01$ in this paper.
As for the Extrusion Label, each extrusion mesh is first built using pythonOCC [24] by the CAD sequence. Then, the SDF of the input points to each extrusion boundary defined by its mesh is calculated. The extrusion label of each point is set as its closest extrusion mesh.

### v3

@@ Abstract
Recovering CAD models from point clouds requires reconstructing their topology and sketch-based extrusion primitives. A dominant paradigm for representing sketches involves implicit neural representations such as Signed Distance Fields (SDFs). However, this indirect approach inherently struggles with precision, leading to unintended curved edges and models that are difficult to edit. In this paper, we propose Point2Primitive, a framework that learns to directly predict the explicit, parametric primitives of CAD models. Our method treats sketch reconstruction as a set prediction problem, employing a improved transformer-based decoder with explicit position queries to directly detect and predict the fundamental sketch curves (i.e., type and parameter) from the point cloud. Instead of approximating a continuous field, we formulate curve parameters as explicit position queries, which are optimized autoregressively to achieve high accuracy. The overall topology is rebuilt via extrusion segmentation. Extensive experiments demonstrate that this direct prediction paradigm significantly outperforms implicit methods in both primitive accuracy and overall geometric fidelity.
@@ 1 Introduction
Computer-Aided Design (CAD) reconstruction, which aims to convert unstructured data like point clouds into structured, parametric models, is a long-standing goal in computer graphics and engineering (WOS:000072777200021). Among various modeling techniques, the sketch-and-extrude process is one of the most fundamental operations, capable of generating a vast array of mechanical parts and industrial designs (willis2021fusion360gallerydataset). Therefore, automatically reversing this process, transforming a raw 3D scan into a structured sketch-based CAD representation, holds immense value for digital manufacturing, legacy part recreation, and design modification.
Despite recent progress driven by large-scale datasets (willis2021fusion360gallerydataset; koch2019abcbigcadmodel), a critical challenge lies in accurately capturing the model’s boundaries. The unstructured, noisy nature of point clouds contrasts sharply with the precise, topologically-defined structure of CAD models. A dominant paradigm in recent literature has been to represent 2D sketches using implicit neural representations such as Signed Distance Fields (SDFs) (9880393; Ren_Zheng_Cai_Li_Zhang_2022; Li_Guo_Zhang_Yan_2023). While these methods are powerful for capturing complex, continuous surfaces, they remain limited in several critical aspects. First, SDFs are purely geometric fields that lack semantic structure, offering no direct correspondence to the primitive entities used in CAD. Second, as approximation-based representations, they inevitably introduce noise and blur fine geometric details, which makes it difficult to recover sharp corners or well-defined curves. Consequently, reconstructed sketches often exhibit smoothed or rounded edges rather than precise boundaries. Finally, because SDF-based sketches cannot be directly expressed in terms of editable primitives—lines, circles, and arcs, the lingua franca of CAD software—an additional conversion step is required. This transformation not only complicates the workflow but also accumulates parameter errors, further hindering faithful reconstruction and downstream editability.
The above limitations lead us to a fundamental question:
Can we bypass the imprecise, intermediate representations and directly predict the explicit, parametric sketch primitives from the point cloud?
In this work, we affirmatively answer this question by proposing Point2Primitive, a framework grounded in a direct prediction methodology for CAD reconstruction from point clouds. Unlike indirect approaches that rely on intermediate surface fitting or voxelization, our method directly targets the underlying parametric sketch primitives. Specifically, we observe that many CAD models can be systematically deconstructed into a sequence of extrusion operations, each defined by a two-dimensional sketch—a closed profile formed by geometric primitives such as lines, arcs, and circles—together with an associated three-dimensional operation that specifies the extrusion direction and length.
Building on this insight, we reformulate the complex reverse engineering problem into two interdependent sub-tasks. First, in the stage of topology reconstruction, the input point cloud is partitioned into clusters that correspond to individual extrusions, a process we term extrusion segmentation. This step disentangles the overall geometry into meaningful sub-structures, enabling subsequent sketch prediction to operate on localized and semantically coherent point sets. Second, in the stage of primitive prediction, we recover the generating two-dimensional sketch for each segmented extrusion body. To achieve this, we design an improved transformer-based decoder that refines point features and directly decodes them into parametric primitives. The decoder integrates explicit geometric priors by formulating curve parameters as positional queries, which guide the attention mechanism and progressively update predictions across layers. This design enhances the accuracy and stability of primitive recovery.
Point2Primitive could bridge topology recovery and parametric sketch inference within a unified direct-prediction framework. By directly predicting the language of CAD, our method produces CAD models that are not only geometrically accurate but also truly editable. In summary, our main contributions are summarized as follows:
  • We introduce Point2Primitive, a new pipeline that directly predicts parametric sketch primitives (i.e., lines, arcs, and circles) from raw point clouds, bypassing implicit fields and preserving sharp, semantically meaningful boundaries for downstream editability.
  • We decompose reconstruction into (i) extrusion segmentation to recover topology by clustering the point cloud into extrusion bodies, and (ii) primitive prediction with an improved transformer decoder that uses explicit position queries and autoregressive parameter refinement to yield high-precision primitive types and parameters.
  • Extensive experiments and ablations show consistent gains in boundary fidelity, primitive recovery, and editability over SDF-based and recent learning baselines, demonstrating both accuracy and robustness across diverse CAD geometries.
@@ 2.1 CAD Reconstruction via Primitive Detection
Several learning-based approaches have also been proposed to fit geometric primitives to point clouds. These methods first segment the input points into clusters corresponding to the same surface primitives and fit the clustered points to the parametric surface primitives (Li_Sung_Dubrovina_Yi_Guibas_2019; Sharma_Liu_Maji; Le_Sung_Ceylan_Mech_Boubekeur_Mitra_2021; RN21) or free surfaces (RN20). Except for surface primitives, UCSG-Net (kania2020ucsgnetunsuperviseddiscovering) and CSG-stump (ren2021csgstumplearningfriendlycsglike) learn the CSG parse tree that interrelates the predicted geometry primitives to reconstruct shapes. The reconstructed CAD models are in B-rep format, which is not as easy to edit as the sketch-and-extrude modeling procedure. In this paper, we explore the inverse of the sketch-and-extrude process and strive to recover the CAD modeling procedure from the point cloud. Therefore, we segment the point clouds into clusters belonging to the same extrusion primitive and explicitly reconstruct every element of the extrusion primitives. Instead of fitting points to the parametric or free surface, we employ an improved transformer to learn and predict the curves directly.
@@ 2.2 Sketch-and-extrude CAD Reconstruction
Some new solutions use SDF to represent sketches following the recent implicit 3D shape representations (Park_Florence_Straub_Newcombe_Lovegrove_2019; wang2022nerf). ExtrudeNet (RN93) and SECAD-Net (Li_Guo_Zhang_Yan_2023) can learn implicit sketches and differentiable extrusions. Point2Cyl reconstructs 3D shapes through sketch regression supervised by the latent embeddings of its SDF. However, these methods utilize the implicit fields for sketch representation, leading to curved edges of the reconstructed shapes. Meanwhile, the reconstructed sketch is hard to be directly edited because it is represented by the SDF. Further transformation from SDF to sketch further brings in extra parameter errors.
@@ 2.3 CAD Generation
Early attempts explored CAD sequence generation with VAE-based models (Wu_Xiao_Zheng_2021; kingma2022autoencodingvariationalbayes) and their extensions such as vector-quantized VAE (xu2023hierarchicalneuralcodingcontrollable) or multimodal diffusion frameworks (MaCLLZ24). However, tokenizing CAD sequences often disrupts the underlying CSG structure (yu2021caprinetlearningcompactcad), making it difficult for language models to capture topology. While SkexGen (Xu_Willis_Lambourne_Cheng_Jayaraman_Furukawa_2022) addressed this by separately encoding topology and geometry, existing VAE–LM hybrids still generalize poorly to reconstruction tasks. Moreover, LLMs struggle with numerical precision (yang2024llmi3dempoweringllm3d), leading to parameter inaccuracies and limited transferability beyond training data.
3 Method
3.1 Preliminaries
A 2D CAD sketch can be modeled as a finite sequence of geometric primitives $S_{i}=(\,c^{(i)}_{1},\ldots,c^{(i)}_{N_{i}}\,)$, where each primitive $c^{(i)}_{j}=(t^{(i)}_{j},\rho^{(i)}_{j})$ comprises a type $t^{(i)}_{j}\in\mathcal{T}$ and a parameter vector $\rho^{(i)}_{j}\in\mathbb{R}^{6}$. We adopt three primitive types $\mathcal{T}=\{\mathsf{L},\mathsf{A},\mathsf{C}\}$ for line, arc, and circle, respectively. Given the point set of the $i$-th extrusion, $\textbf{P}_{\mathcal{E}_{i}}\in\mathbb{R}^{N_{\mathcal{E}_{i}}\times 3}$, the network $f_{\theta}$ is trained to predict the sketch as a set/sequence of primitives:
$f_{\theta}\!\left(\{\,\mathbf{p}_{k}\mid\mathbf{p}_{k}\!\in\!\textbf{P}_{\mathcal{E}_{i}}\,\}\right)\;\approx\;S_{i}\;=\;(\,c^{(i)}_{1},\ldots,c^{(i)}_{N_{i}}\,),\quad c^{(i)}_{j}=(t^{(i)}_{j},\rho^{(i)}_{j}),\;t^{(i)}_{j}\!\in\!\{\mathsf{L},\mathsf{A},\mathsf{C}\}.$ (1)
Center-prior parameterization.
We use a center-prior parameterization with a fixed 6-D vector $\rho=(\rho_{1},\ldots,\rho_{6})$ and pad unused entries by 1-1. Line ($\mathsf{L}$): midpoint $\mathbf{m}=(x_{m},y_{m})$ and one endpoint $\mathbf{d}_{1}=(x_{1},y_{1})$, encoded as $\rho=(x_{m},y_{m},x_{1},y_{1},-1,-1)$. Arc ($\mathsf{A}$): midpoint $\mathbf{m}=(x_{m},y_{m})$, start point $\mathbf{s}=(x_{1},y_{1})$, and end point $\mathbf{d}_{2}=(x_{2},y_{2})$, encoded as $\rho=(x_{m},y_{m},x_{1},y_{1},x_{2},y_{2})$. Circle ($\mathsf{C}$): center $\mathbf{c}=(x_{c},y_{c})$ and radius $r$, encoded as $\rho=(x_{c},y_{c},r,-1,-1,-1)$. Unless otherwise stated, we refer to a primitive by $c=(t,\rho)$. The first two entries of $\rho$ (i.e., the center or midpoint) are further used as positional embeddings to guide attention in the improved transformer decoder (see Sec. 3.4).
Sketch–extrude–shape relation.
An extrusion is denoted by $\mathcal{E}_{i}=[S_{i},E_{i}]$, where $E_{i}$ is the associated operation. A solid shape is a finite collection of extrusions, $\mathbb{S}=\{\mathcal{E}_{1},\ldots,\mathcal{E}_{K}\}$, and is represented by the signed distance field (SDF) induced by these extrusions. In our pipeline, $\{\mathcal{E}_{1},\ldots,\mathcal{E}_{i}\}$ is recovered via point segmentation, while the current sketch $S_{i}$ is predicted with the improved transformer decoder.
Based on the above preliminaries, we present Point2Primitive, a direct sketch primitives prediction method for CAD reconstruction from point clouds. As shown in Figure 1, our method first segments points into extrusions, classifies base/barrel points, and then applies an improved transformer to predict curve types and parameters as a set-to-set problem. The segmentation head builds on PointCNN with fixed grid points, while the transformer refines features and enhances prediction accuracy.
@@ 3.2 CAD Dataset Generation
In this paper, we propose to generate the training dataset based on the Signed Distance Field (SDF). An SDF $\texttt{SDF}(p_{\text{eval}},b)$ is a continuous function that, for a given spatial eval point $p_{\text{eval}_{i}}$, outputs the eval point’s distance to the closest boundary defined by $b$, whose sign denotes whether the point is inside (negative) or outside (positive) of the watertight surface. The data generation pipeline can convert any CAD sequence saved in JSON files following the style of the Fusion360 Gallery dataset (willis2021fusion360gallerydataset) to the training data pairs.
Given an extrusion point set $\textbf{P}_{\mathcal{E}_{i}}\in R^{N_{\mathcal{E}_{i}}\times{3}}$ and its sketch $S_{i}$. We first project the points $\textbf{P}_{\mathcal{E}_{i}}$ to the sketch plane as eval points according to the transform matrix $W_{i}$ defined by its extrusion axis $e_{i}$. Then, its SDF is calculated to the boundary defined by the sketch $S_{i}$. The batch SDF calculation will be detailed in the Appendix F.1. The barrel label $b_{i}$ of each point can be defined as follows:
$b_{i}=\texttt{Bool}(\texttt{SDF}(W_{i}\cdot{\textbf{P}_{\mathcal{E}_{i}}},S_{i})\leq{\rm thresh})\,$ (2)
where $\rm thresh$} controls the label precision and ${\rm thresh}=0.01$ in this paper.
As for the Extrusion Label, each extrusion mesh is first built using pythonOCC (thomas_paviot_2022_7471333) by the CAD sequence. Then, the SDF of the input points to each extrusion boundary defined by its mesh is calculated. The extrusion label of each point is set as its closest extrusion mesh.